# CQ-DINO: Mitigating Gradient Dilution via Category Queries for Vast Vocabulary Object Detection

### v3

@@ 3.4 Encoding Category Correlations
Explicit Hierarchical Tree Construction. For hierarchical datasets like V3Det [35], we introduce hierarchical tree construction, as shown in Fig. 5. The process begins at leaf nodes and progressively integrates hierarchical information upward through the category tree. For each parent node $v$ with children $\mathcal{C}(v)$, its new query $Q_{v}^{\prime}$ is a combination of its original query $Q_{v}$ and the mean pooling of all direct child nodes. Leaf nodes retain their original features directly, since they have no children.
@@ 4.1 Datasets and Implementation Details
We conduct experiments on two detection benchmarks: (1) V3Det [35]: a vast vocabulary detection dataset containing 13,204 categories, with 183k training and 30k validation images. This is our primary benchmark for evaluating vast vocabulary detection. (2) COCO val2017 [19]: a standard benchmark dataset with 80 object categories, comprising 118k training and 5k validation images. We include this dataset to verify the effectiveness of our method in limited vocabulary scenarios.
@@
Method Epochs Backbone $AP$ $AP_{50}$ $AP_{75}$
ATSS  [42] 24 Swin-B 7.6 8.9 8.0
FCOS [33] 24 Swin-B 21.0 24.8 22.3
Faster R-CNN [27] 24 Swin-B 37.6 46.0 41.1
CenterNet2 [46] 24 Swin-B 39.8 46.1 42.4
Cascade R-CNN [2] 24 Swin-B 42.5 49.1 44.9
Deformable DETR [47] 50 Swin-B 42.5 48.3 44.7
DINO [41] 24 Swin-B 42.0 46.8 43.9
Prova [4] 24 Swin-B 44.5 49.9 46.6
@@
DINO [41] 24 Swin-B-22k 43.4 48.4 45.4
Prova [4] 24 Swin-B-22k 50.3 56.1 52.6
@@
DINO [41] 24 Swin-L 48.5 54.3 50.7
Prova [4] 24 Swin-L 50.9 57.2 53.2
@@
Method Epochs Backbone $AP$ $AP_{S}$ $AP_{M}$ $AP_{L}$ H-Def-DETR [13] 36 Swin-L 57.1 39.7 61.4 73.4 Relation-DETR [12] 12 Swin-L 57.8 41.2 62.1 74.4 DINO [41] 36 Swin-L 58.0 41.3 62.1 73.6 Rank-DETR [25] 36 Swin-L 58.2 42.4 62.2 73.6 CQ-DINO (ours) 24 Swin-L 58.5 42.5 62.1 74.0
@@
Method Backbone $AP$ $AP_{50}$ $AP_{75}$ GenerateU [18] Swin-L&T5-B 0.4 0.5 0.4 ChatRex [14] Swin-L&LLM-7B 1.3 1.5 1.4 GenerateU* [18] Swin-L&T5-B 21.8 27.2 22.1 DINO [41] Swin-L 48.5 54.4 50.7 CQ-DINO (Ours) Swin-L 53.0 58.4 55.4
Experiments are conducted on 8 A100-40G GPUs with a total batch size of 16, unless otherwise specified. Baseline configurations are used for fair comparison. Three Swin Transformer [23] variants serve as backbones: Swin-B (ImageNet-1k [7] pre-trained), Swin-B-22k (ImageNet-22k pre-trained), and Swin-L (ImageNet-22k pre-trained). Category queries are initialized using CLIP-ViT-L [26] text embeddings. To match dataset vocabularies, we employ 100 category queries for V3Det and 30 for COCO, aligned with their category sizes. The hierarchical tree structure is derived from V3Det’s category taxonomy. For implicit relation learning, we use an 8-head self-attention module.
The training objective combines multiple loss terms with the following weights: classification loss (Asymmetric Loss [29], weight=1.0), contrastive alignment (Focal Loss [30], weight=1.0), bounding box regression (L1 Loss, weight=5.0), and GIoU Loss [28] (weight=2.0), as in GroundingDINO [22]. Hungarian matching is used, following GroundingDINO, with identical matching costs for object-to-query assignment. To stabilize training, we use a two-stage approach: first, pre-training category queries, image encoder, and image-guided query selection for 10 epochs to establish high initial target category recall; then, fine-tuning the full detection pipeline in the second stage.
@@ 4.2 Experimental Results
Performance on Vast Vocabulary Detection. Tab. 1 presents our comparison with state-of-the-art methods on the V3Det [35] benchmark. CQ-DINO consistently outperforms all previous approaches across different backbone configurations. With the Swin-B backbone, CQ-DINO achieves 46.3% AP, outperforming general detection methods like Deformable DETR [47] by 3.8% AP and DINO [41] by 4.3% AP. More importantly, CQ-DINO surpasses Prova [4], a specialized vast vocabulary detection method, by 1.8% AP. When integrated with the Swin-B-22k backbone, CQ-DINO achieves 52.3% AP, outperforming Prova by 2.0% AP. With the Swin-L backbone, CQ-DINO achieves 53.0% AP. The consistent improvements across different backbones demonstrate that CQ-DINO effectively addresses vast vocabulary detection challenges.
Performance on Standard Detection Benchmark. Tab. 3 compares our method with state-of-the-art approaches on COCO val2017. Despite being primarily designed for vast vocabulary scenarios, CQ-DINO achieves competitive performance, reaching 58.5% AP, which is comparable to recent DETR-based methods. We report the best results from their original papers to ensure a fair comparison. The competitive performance of CQ-DINO is mainly due to the proposed gradient rebalancing and adaptive hard mining strategies.
@@
Encoding Category Correlations Image-guided Query Selection $AP$ $AR^{C}$ FPS
@@ 4.3 Ablation Studies
Effect of Each Component in CQ-DINO. Table 4 presents the contribution of each component in terms of average precision ($AP$) and category-level average recall with selected queries ($AR^{C}$). To establish a baseline without image-guided query selection, we conduct experiments on 8 H800-80G GPUs due to memory constraints. Notably, incorporating image-guided query selection increases FPS from 0.7 to 10.8, highlighting its effectiveness in alleviating memory bottlenecks and substantially enhancing inference efficiency. Furthermore, this component addresses the issue of gradient dilution and leads to a substantial AP improvement, from 47.3% to 51.1% in detection performance.
@@
Method $AP$ $AR^{C}$ Params (M) CQ-DINO w/o SA 58.3 98.2 244.3 CQ-DINO w/ SA 58.5 ($\uparrow 0.2$) 99.1 ($\uparrow 0.9$) 246.7 ($+2.4$)
@@
Method $AP$ $AP_{50}$ $AP_{75}$ $AR^{C}$ Fixed weight (0.5) 51.9 57.4 54.3 82.3 Ours ($\alpha_{v}$) 52.3 57.7 55.4 83.3
Effectiveness of Adaptive Weighting in Tree Construction. Tab. 6 shows the importance of our adaptive weighting strategy compared to a fixed weight of 0.5. This approach adjusts weights based on the varying number of child categories for each parent node in the hierarchy. Results show that our adaptive approach outperforms fixed weighting.
@@ 4.4 Discussion
Method Params/Cat. Memory/Cat. Max Cats. (K) (kB) (k) DINO [41] 2.1 8.9 100 CQ-DINO 0.8 2.7 130
@@
Limitations of Generation-based Methods. We evaluate generation-based methods on the V3Det dataset in Tab. 3. Following the evaluation protocol from GenerateU [18], we compute semantic similarity between generated category embeddings and V3Det category embeddings, selecting the highest similarity match as the final prediction. Our experiments reveal poor performance: GenerateU [18] achieves only 0.4% AP, while the more recent ChatRex [14] achieves just 1.3% AP. This highlights a fundamental limitation: generation-based methods struggle to control the granularity of generated categories, creating semantic misalignments with the specific requirements of detection tasks. Furthermore, even when we finetune GenerateU on V3Det data, the resulting performance (21.8% AP) still exhibits a substantial gap compared to classification-head methods.
Focal Loss Parameters Analysis under Gradient Dilution. As discussed in Sec. 3.1, vast vocabulary detection suffers from gradient dilution challenges. Focal Loss (FL) [30] mitigates this via adaptive weighting with hyperparameters $\alpha$ and $\gamma$. Theoretically, $\alpha$ balances positive/negative sample contributions, and increasing $\alpha$ enhances the model’s ability to handle more categories. The factor $\gamma$ focuses learning on hard negatives, where increasing $\gamma$ improves the mining of hard examples. Tab. 8 evaluates FL configurations on DINO. The default setting [35] ($\alpha=0.25$, $\gamma=2$) achieves 43.4% AP. Tuning these hyperparameters reveals that training is unstable when $\gamma\geq 5$ or $\alpha\geq 0.75$. Notably, the optimal configuration ($\gamma=2$ and $\alpha=0.5$) achieves 47.4% AP, surpassing the baseline by 4.0%. Nonetheless, this is still lower than the 52.3% AP achieved by our CQ-DINO, suggesting that FL hyperparameter tuning, while beneficial, leaves room for further improvement. Interestingly, when examining parameter transferability across architectures in Appendix Tab. 9, we find that this optimal setting does not generalize well. Applying the Swin-B-22k optimal parameters ($\alpha=0.5$, $\gamma=2$) to Swin-B degrades performance by 3.3 % AP relative to its default setting. However, they increase the performance by 4.0% AP and 1.6% AP for Swin-B-22k and Swin-L backbones, respectively. These findings suggest that while Focal Loss is effective in addressing the gradient dilution challenge, optimal hyperparameter selection remains architecture-dependent and requires careful tuning.
@@ Appendix A Appendix
A.1 Performance of DINO with Different Focal Loss Parameters

### v4

@@ 3.4 Encoding Category Correlations
Explicit Hierarchical Tree Construction.  For hierarchical datasets like V3Det wang2023v3det , we introduce hierarchical tree construction, as shown in Fig. 5. The process begins at leaf nodes and progressively integrates hierarchical information upward through the category tree. For each parent node $v$ with children $\mathcal{C}(v)$, its new query $Q_{v}^{\prime}$ is a combination of its original query $Q_{v}$ and the mean pooling of all direct child nodes. Leaf nodes retain their original features directly, since they have no children.
@@ 4.1 Datasets and Implementation Details
We conduct experiments on two detection benchmarks: (1) V3Det wang2023v3det : a vast vocabulary detection dataset containing 13,204 categories, with 183k training and 30k validation images. This is our primary benchmark for evaluating vast vocabulary detection. (2) COCO val2017 coco : a standard benchmark dataset with 80 object categories, comprising 118k training and 5k validation images. We include this dataset to verify the effectiveness of our method in limited vocabulary scenarios.
@@
Method Epochs Backbone APAP $AP_{50}$ $AP_{75}$
ATSS  atss 24 Swin-B 7.6 8.9 8.0
FCOS fcos 24 Swin-B 21.0 24.8 22.3
Faster R-CNN fasterrcnn 24 Swin-B 37.6 46.0 41.1
CenterNet2 centernet2 24 Swin-B 39.8 46.1 42.4
Cascade R-CNN cascadercnn 24 Swin-B 42.5 49.1 44.9
Deformable DETR zhu2021deformable 50 Swin-B 42.5 48.3 44.7
DINO zhang2023dino 24 Swin-B 42.0 46.8 43.9
Prova chen2024comprehensive 24 Swin-B 44.5 49.9 46.6
@@
DINO zhang2023dino 24 Swin-B-22k 43.4 48.4 45.4
Prova chen2024comprehensive 24 Swin-B-22k 50.3 56.1 52.6
@@
DINO zhang2023dino 24 Swin-L 48.5 54.3 50.7
Prova chen2024comprehensive 24 Swin-L 50.9 57.2 53.2
@@
Method Epochs Backbone APAP $AP_{S}$ $AP_{M}$ $AP_{L}$ H-Def-DETR hfdetr 36 Swin-L 57.1 39.7 61.4 73.4 Relation-DETR hou2024relation 12 Swin-L 57.8 41.2 62.1 74.4 DINO zhang2023dino 36 Swin-L 58.0 41.3 62.1 73.6 Rank-DETR pu2023rank 36 Swin-L 58.2 42.4 62.2 73.6 CQ-DINO (ours) 24 Swin-L 58.5 42.5 62.1 74.0
@@
Method Backbone APAP $AP_{50}$ $AP_{75}$ GenerateU generateu Swin-L&T5-B 0.4 0.5 0.4 ChatRex jiang2024chatrex Swin-L&LLM-7B 1.3 1.5 1.4 GenerateU* generateu Swin-L&T5-B 21.8 27.2 22.1 DINO zhang2023dino Swin-L 48.5 54.4 50.7 CQ-DINO (Ours) Swin-L 53.0 58.4 55.4
Experiments are conducted on 8 A100-40G GPUs with a total batch size of 16, unless otherwise specified. Baseline configurations are used for fair comparison. Three Swin Transformer liu2021Swin variants serve as backbones: Swin-B (ImageNet-1k deng2009imagenet pre-trained), Swin-B-22k (ImageNet-22k pre-trained), and Swin-L (ImageNet-22k pre-trained). Category queries are initialized using CLIP-ViT-L clip text embeddings. To match dataset vocabularies, we employ 100 category queries for V3Det and 30 for COCO, aligned with their category sizes. The hierarchical tree structure is derived from V3Det’s category taxonomy. For implicit relation learning, we use an 8-head self-attention module.
The training objective combines multiple loss terms with the following weights: classification loss (Asymmetric Loss ridnik2021asymmetric , weight=1.0), contrastive alignment (Focal Loss ross2017focal , weight=1.0), bounding box regression (L1 Loss, weight=5.0), and GIoU Loss giou (weight=2.0), as in GroundingDINO gdino . Hungarian matching is used, following GroundingDINO, with identical matching costs for object-to-query assignment. To stabilize training, we use a two-stage approach: first, pre-training category queries, image encoder, and image-guided query selection for 10 epochs to establish high initial target category recall; then, fine-tuning the full detection pipeline in the second stage.
@@ 4.2 Experimental Results
Performance on Vast Vocabulary Detection. Tab. 1 presents our comparison with state-of-the-art methods on the V3Det wang2023v3det benchmark. CQ-DINO consistently outperforms all previous approaches across different backbone configurations. With the Swin-B backbone, CQ-DINO achieves 46.3% AP, outperforming general detection methods like Deformable DETR zhu2021deformable by 3.8% AP and DINO zhang2023dino by 4.3% AP. More importantly, CQ-DINO surpasses Prova chen2024comprehensive , a specialized vast vocabulary detection method, by 1.8% AP. When integrated with the Swin-B-22k backbone, CQ-DINO achieves 52.3% AP, outperforming Prova by 2.0% AP. With the Swin-L backbone, CQ-DINO achieves 53.0% AP. The consistent improvements across different backbones demonstrate that CQ-DINO effectively addresses vast vocabulary detection challenges.
Performance on Standard Detection Benchmark
. Tab. 3 compares our method with state-of-the-art approaches on COCO val2017. Despite being primarily designed for vast vocabulary scenarios, CQ-DINO achieves competitive performance, reaching 58.5% AP, which is comparable to recent DETR-based methods. We report the best results from their original papers to ensure a fair comparison. The competitive performance of CQ-DINO is mainly due to the proposed gradient rebalancing and adaptive hard mining strategies.
@@
Encoding Category Correlations Image-guided Query Selection APAP $AR^{C}$ FPS
@@ 4.3 Ablation Studies
Effect of Each Component in CQ-DINO.
Table 4 presents the contribution of each component in terms of average precision ($AP$) and category-level average recall with selected queries ($AR^{C}$). To establish a baseline without image-guided query selection, we conduct experiments on 8 H800-80G GPUs due to memory constraints. Notably, incorporating image-guided query selection increases FPS from 0.7 to 10.8, highlighting its effectiveness in alleviating memory bottlenecks and substantially enhancing inference efficiency. Furthermore, this component addresses the issue of gradient dilution and leads to a substantial AP improvement, from 47.3% to 51.1% in detection performance.
@@
Method APAP $AR^{C}$ Params (M) CQ-DINO w/o SA 58.3 98.2 244.3 CQ-DINO w/ SA 58.5 ($\uparrow 0.2$) 99.1 ($\uparrow 0.9$) 246.7 (+2.4+2.4)
@@
Method APAP $AP_{50}$ $AP_{75}$ $AR^{C}$ Fixed weight (0.5) 51.9 57.4 54.3 82.3 Ours ($\alpha_{v}$) 52.3 57.7 55.4 83.3
Effectiveness of Adaptive Weighting in Tree Construction.
Tab. 6 shows the importance of our adaptive weighting strategy compared to a fixed weight of 0.5. This approach adjusts weights based on the varying number of child categories for each parent node in the hierarchy. Results show that our adaptive approach outperforms fixed weighting.
@@ 4.4 Discussion
Method Params/Cat. Memory/Cat. Max Cats. (K) (kB) (k) DINO zhang2023dino 2.1 8.9 100 CQ-DINO 0.8 2.7 130
@@
Limitations of Generation-based Methods.
We evaluate generation-based methods on the V3Det dataset in Tab. 3. Following the evaluation protocol from GenerateU generateu , we compute semantic similarity between generated category embeddings and V3Det category embeddings, selecting the highest similarity match as the final prediction. Our experiments reveal poor performance: GenerateU generateu achieves only 0.4% AP, while the more recent ChatRex jiang2024chatrex achieves just 1.3% AP. This highlights a fundamental limitation: generation-based methods struggle to control the granularity of generated categories, creating semantic misalignments with the specific requirements of detection tasks. Furthermore, even when we finetune GenerateU on V3Det data, the resulting performance (21.8% AP) still exhibits a substantial gap compared to classification-head methods.
Focal Loss Parameters Analysis under Gradient Dilution.
As discussed in Sec. 3.1, vast vocabulary detection suffers from gradient dilution challenges. Focal Loss (FL) ross2017focal mitigates this via adaptive weighting with hyperparameters $\alpha$ and $\gamma$. Theoretically, $\alpha$ balances positive/negative sample contributions, and increasing $\alpha$ enhances the model’s ability to handle more categories. The factor $\gamma$ focuses learning on hard negatives, where increasing $\gamma$ improves the mining of hard examples. Tab. 8 evaluates FL configurations on DINO. The default setting wang2023v3det ($\alpha=0.25$, $\gamma=2$) achieves 43.4% AP. Tuning these hyperparameters reveals that training is unstable when $\gamma\geq 5$ or $\alpha\geq 0.75$. Notably, the optimal configuration ($\gamma=2$ and $\alpha=0.5$) achieves 47.4% AP, surpassing the baseline by 4.0%. Nonetheless, this is still lower than the 52.3% AP achieved by our CQ-DINO, suggesting that FL hyperparameter tuning, while beneficial, leaves room for further improvement. Interestingly, when examining parameter transferability across architectures in Appendix Tab. 9, we find that this optimal setting does not generalize well. Applying the Swin-B-22k optimal parameters ($\alpha=0.5$, $\gamma=2$) to Swin-B degrades performance by 3.3 % AP relative to its default setting. However, they increase the performance by 4.0% AP and 1.6% AP for Swin-B-22k and Swin-L backbones, respectively. These findings suggest that while Focal Loss is effective in addressing the gradient dilution challenge, optimal hyperparameter selection remains architecture-dependent and requires careful tuning.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by NSFC under Grant 62222112 and 62176186; and in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017.
@@ Appendix A Appendix
A.1 Extended Analysis on Gradient Dilution in Vast Vocabulary vs. Class-Imbalanced
In Sec. 3.1 of the main paper, we introduced the concept of positive gradient dilution as a primary challenge in vast vocabulary object detection, using a simplified model assuming a relatively balanced data distribution. This appendix provides a more comprehensive analysis to further clarify the distinction between the gradient dilution problem caused by vast vocabulary size and the one addressed in traditional class-imbalanced learning.
The simplified gradient ratio in our main paper (Eq. (3)) demonstrates that $\rho\propto\frac{1}{C}$, highlighting the direct impact of the vocabulary size $C$. To better distinguish the effects of class imbalance and vocabulary size, we can formulate a more general gradient signal ratio, $\rho_{c^{+}}$, for a positive class $c^{+}$:
$\rho_{c^{+}}=\frac{n_{c^{+}}\cdot\|\nabla_{z_{c^{+}}}\mathcal{L}\|}{\sum_{c^{-}\neq c^{+}}^{C}n_{c^{-}}\cdot\|\nabla_{z_{c^{-}}}\mathcal{L}\|}\approx\frac{n_{c^{+}}\cdot(1-\sigma(z_{c^{+}}))}{\mathbb{E}_{c^{-}}[\sigma(z_{c^{-}})]\cdot\sum_{c^{-}\neq c^{+}}^{C}n_{c^{-}}}$
By letting $\epsilon^{+}$ and $\epsilon^{-}$ represent the average gradient magnitudes for positive and negative samples respectively, and $N$ be the total sample count across all classes, the ratio can be expressed as:
$\rho_{c^{+}}\propto\frac{n_{c^{+}}\cdot\epsilon^{+}}{(N-n_{c^{+}})\cdot\epsilon^{-}}$ (8)
This generalized formula reveals two distinct and compounding sources of gradient dilution:
Class Imbalance: This well-studied issue is primarily reflected by the term $n_{c^{+}}$ in the numerator. When a class is rare (i.e., has a long-tailed distribution), its sample count $n_{c^{+}}$ is small, which directly reduces the gradient signal ratio $\rho_{c^{+}}$. This is the central challenge that traditional class-imbalance methods aim to solve.
Vast Vocabulary Size: This is primarily driven by the term $N-n_{c^{+}}$ in the denominator. In vast vocabulary settings ($C>10,000$), $N-n_{c^{+}}$ becomes enormous because it aggregates all negative samples from the other C1C-1 categories. The large $C$ further reduces the ratio $\rho_{c^{+}}$, making the gradient dilution more severe in vast vocabulary object detection. This impacts both rare and common classes.
This distinction in the problem’s source explains why conventional methods for class imbalance are not sufficient for the vast vocabulary challenge.
A.2 Algorithm Details
This section provides the detailed algorithmic implementations of the three core components in CQ-DINO, as described in Sec. 3. Each algorithm addresses specific challenges in vast vocabulary object detection:
Algorithm 1 - Image-Guided Query Selection (Sec. 3.3): This algorithm implements the core innovation of CQ-DINO by dynamically selecting the most relevant category queries for each image. Through cross-attention mechanisms between category queries and image features, it reduces the negative search space from the full vocabulary to a manageable subset ($C^{\prime}\ll C$), addressing both positive gradient dilution and hard negative gradient dilution issues identified in Sec. 3.1.
Input:
  $B$: Batch size, $C$: Number of category queries,
  $D$: Embedding dimension, $C^{\prime}$: Number of selected category queries,
  $H^{\prime}$: Height of image features, $W^{\prime}$: Width of image features,
  $\mathbf{Q}_{cat}\in\mathbb{R}^{B\times C\times D}$: Category queries,
  $\mathbf{F}_{img}\in\mathbb{R}^{B\times D\times H^{\prime}\times W^{\prime}}$: Image features
Output:
  $\mathbf{Q}_{cat}^{\prime}\in\mathbb{R}^{B\times C^{\prime}\times D}$: Selected enhanced category queries,
  $\mathcal{I}\in\mathbb{R}^{B\times C^{\prime}}$: Selection indices
// Reshape image features for cross-attention
$\mathbf{F}_{flat}\leftarrow\text{Reshape}(\mathbf{F}_{img},[B,D,H^{\prime}\times W^{\prime}])$;
$\mathbf{Q}_{enhanced}\leftarrow\mathbf{Q}_{cat}$;
// Enhancement through cross-attention layers
for $l=1$ to 22 do
  // Cross-attention with image features
 $\mathbf{Q}_{attn}\leftarrow\text{MultiHeadCrossAttention}(\mathbf{Q}_{enhanced},\mathbf{F}_{flat},\mathbf{F}_{flat})$;
 $\mathbf{Q}_{enhanced}\leftarrow\text{LayerNorm}(\mathbf{Q}_{attn}+\mathbf{Q}_{enhanced})$;
 // Feed-forward transformation
 $\mathbf{Q}_{ffn}\leftarrow\text{FFN}(\mathbf{Q}_{enhanced})$;
 $\mathbf{Q}_{enhanced}\leftarrow\text{LayerNorm}(\mathbf{Q}_{ffn}+\mathbf{Q}_{enhanced})$;
// Query selection based on enhanced representations
$\mathbf{L}\leftarrow\text{LinearProjection}(\mathbf{Q}_{enhanced})$ // $\mathbf{L}\in\mathbb{R}^{B\times C}$
for $b=1$ to $B$ do
 $\mathcal{I}_{b,:}\leftarrow\text{TopK}(\mathbf{L}_{b,:},C^{\prime})$;
 $\mathbf{Q}_{cat}^{\prime}[b,:,:]\leftarrow\mathbf{Q}_{enhanced}[b,\mathcal{I}_{b,:},:]$;
return $\mathbf{Q}_{cat}^{\prime}$, $\mathcal{I}$
Algorithm 1 Image-Guided Query Selection
Algorithm 2 - Self-Attention for Implicit Category Relations (Sec. 3.4): For datasets without explicit hierarchical structures, this algorithm employs multi-head self-attention to learn implicit correlations between categories. It allows semantically related categories to influence each other’s representations based on learned attention patterns, complementing the explicit hierarchical approach.
Input: $\mathbf{Q}_{cat}\in\mathbb{R}^{C\times D}$: Category queries,
$H$: Number of attention heads
Output: $\mathbf{Q}_{corr}\in\mathbb{R}^{C\times D}$: Correlation-enhanced queries
// Enhance queries through self-attention mechanism
$\mathbf{Q}_{attn}\leftarrow\text{MultiHeadSelfAttention}(\mathbf{Q}_{cat})$;
$\mathbf{Q}_{corr}\leftarrow\text{LayerNorm}(\mathbf{Q}_{attn}+\mathbf{Q}_{cat})$;
return $\mathbf{Q}_{corr}$
Algorithm 2 Self-Attention for Implicit Category Relations
Algorithm 3 - Explicit Hierarchical Tree Construction (Sec. 3.4): For datasets with explicit hierarchical structures like V3Det, this algorithm leverages the category hierarchical structures to enhance query representations. It performs bottom-up tree traversal to incorporate hierarchical relationships through adaptive weighting, enabling parent categories to aggregate semantic information from their children while maintaining individual semantics.
Input: $\mathbf{Q}_{cat}\in\mathbb{R}^{C\times D}$: Category queries,
$\mathcal{T}$: Hierarchical tree structure with nodes $\mathcal{V}$,
$w$: Base weight parameter (default: 0.3)
Output: $\mathbf{Q}_{corr}\in\mathbb{R}^{C\times D}$: Correlation-enhanced queries
// Initialize and prepare tree traversal
$\mathbf{Q}_{corr}\leftarrow\mathbf{Q}_{cat}$;
$N_{max}\leftarrow\max_{v\in\mathcal{V}}|\text{Children}(v,\mathcal{T})|$;
$\mathcal{L}\leftarrow\text{TopologicalSort}(\mathcal{T})$;
// Bottom-up tree traversal for correlation enhancement
foreach node $v$ in $\text{Reverse}(\mathcal{L})$ do
 if $\text{IsLeaf}(v,\mathcal{T})$ then
     // Leaf nodes retain original queries
    continue;
 else
     // Compute adaptive weight based on children count
    $\mathcal{C}(v)\leftarrow\text{GetChildren}(v,\mathcal{T})$;
    $n_{v}\leftarrow|\mathcal{C}(v)|$;
    $\alpha_{v}\leftarrow w+\frac{\log(n_{v}+1)}{\log(N_{max}+1)}$;
    $\alpha_{v}\leftarrow\min(\alpha_{v},1.0)$;
    // Update parent query with weighted combination
    $\mathbf{Q}_{child}^{mean}\leftarrow\frac{1}{n_{v}}\sum_{c\in\mathcal{C}(v)}\mathbf{Q}_{corr}[c,:]$;
    $\mathbf{Q}_{corr}[v,:]\leftarrow(1-\alpha_{v})\cdot\mathbf{Q}_{cat}[v,:]+\alpha_{v}\cdot\mathbf{Q}_{child}^{mean}$;
return $\mathbf{Q}_{corr}$
Algorithm 3 Explicit Hierarchical Tree Construction
A.3 Performance of DINO with Different Focal Loss Parameters